# Injecting Domain-Specific Knowledge into Large Language Models: A Comprehensive Survey

### v1

@@ 1 Introduction
Large Language Models (LLMs) have achieved extraordinary success across various tasks, showcasing remarkable capabilities in reasoning, knowledge representation, and decision-making. However, despite their impressive performance in general-purpose applications, many specialized domains, such as healthcare, chemistry, and legal analysis, demand the integration of domain-specific knowledge to achieve high accuracy and reliability. To address this challenge, researchers have explored methods to enhance LLMs through external or embedded domain expertise, a process often referred to as knowledge injection. This approach aims to bridge the gap between general-purpose language understanding and the stringent requirements of domain-specific tasks, enabling LLMs to perform effectively in highly specialized contexts.
Figure 1: Illustration of Growth Trends in Domain-Specific Knowledge Injection into LLMs. The chart displays the cumulative number of papers published between October 2022 and December 2024. Different colors and border styles represent various injection methods and domains, such as blue with a solid border denoting dynamic injection in the biomedical field.
Building on the foundational capabilities of general-purpose LLMs, knowledge injection techniques provide an effective means to address their limitations in handling specialized applications. Compared to the generalized approach of standard LLMs, knowledge injection offers two key advantages: 1) incorporating precise, domain-specific knowledge to improve accuracy and reliability in specialized tasks, and 2) allowing LLMs to dynamically adapt to new information or evolving knowledge bases, ensuring up-to-date expertise. These techniques bridge the gap between general-purpose understanding and domain-specific demands by leveraging both structured and unstructured knowledge sources. As a result, knowledge injection methods have been successfully applied in fields such as healthcare, chemistry, and legal analysis, significantly enhancing LLM performance. For example, biomedical LLMs  (Bolton et al., 2024; Yan et al., 2023) have demonstrated superior accuracy in tasks like medical diagnostics and regulatory compliance, while domain-specific models for material science (Xie et al., 2024; Antunes et al., 2024; Zhang et al., 2024a) have achieved advances in material property prediction and discovery. These dedicated models underscore the transformative potential of integrating domain knowledge into LLMs, unlocking solutions to complex, field-specific challenges.
@@
To assist readers from various backgrounds in understanding knowledge injection techniques for LLMs and to complement existing surveys by addressing unresolved questions, we organize our survey paper as follows. After providing the foundational background knowledge in Section 2, we address a pivotal question: How can domain-specific knowledge be effectively integrated into LLMs? To answer this, we present a comprehensive framework for categorizing knowledge injection methods in Section 3. We delve into this topic by discussing: 1) Dynamic Knowledge Injection, which explains how external knowledge is retrieved and incorporated in real-time during inference to enhance reasoning capabilities; 2) Static Knowledge Embedding, which describes how domain knowledge is embedded into the model during training or fine-tuning to make it an inherent part of the model; 3) Adapters, which highlight modular techniques for storing and utilizing external knowledge without altering the primary model’s parameters; and 4) Prompt Optimization, which focuses on how carefully designed prompts enable models to utilize existing knowledge without requiring changes to their architecture. Another perspective for reviewing knowledge injection studies is their application across various domains. In Section 4, we categorize current application domains, such as materials science, chemistry, biology, and law. To guide the identification of appropriate tools and resources, we summarize commonly used benchmarks, open-source frameworks, and analyses on performance in Section 5. Based on these summaries, we discuss the challenges and opportunities in this evolving field in Section 6. The conclusions are summarized in Section 7.
@@ 2.1 Domain-Specific Knowledge
Domain-specific knowledge refers to specialized information or expertise pertinent to a specific field or application, distinguishing it from general knowledge that spans across multiple domains. While general knowledge enables models to understand broad contexts, domain-specific knowledge is essential for addressing specialized tasks where precise, field-specific understanding is required. For instance, in scientific text processing (Bran et al., 2023), models must comprehend complex scientific terminologies, concepts, and methodologies to provide accurate and relevant answers. Similarly, in e-commerce search (Zhao et al., 2024a), understanding domain-specific terms such as product categories, technical specifications, or colloquial shopping language is crucial for delivering relevant search results and recommendations. In healthcare applications, LLMs must understand medical terminologies, diagnoses, treatment plans, and drug interactions. For example, biomedical question answering (Pei and others, 2024) and medical report summarization rely on integrating knowledge from medical literature like PubMed (Dernoncourt and Lee, 2017).
To address these needs, researchers have explored various methods for incorporating domain-specific knowledge into LLMs. In this paper, we aim to provide a survey of these various injection methods.
@@ 2.2 Knowledge Representation and Encoding
Knowledge can take different forms depending on the structure and application needs. For example, knowledge graphs (Zhang et al., 2024c) represent information as entities and relationships in a graph format, allowing structured reasoning and inference. These are widely used in tasks like question-answering and recommendation systems, where relationships between entities are crucial. Similarly, knowledge in text form, such as Wikipedia (Jeong et al., 2024), provides a vast corpus of unstructured information.
Knowledge can also be stored in vector space instead of readable text or graph formats. For instance, soft prompt tuning (Singhal et al., 2023a) learns useful knowledge in vector form, which is concatenated with the original input to guide LLMs in performing specific downstream tasks. In addition to external representations, knowledge can also emerge from within the model itself. For example, chain-of-thought prompting (Yao et al., 2024) introduces intermediate reasoning steps that help the model break down complex tasks into manageable parts. By explicitly reasoning through these steps, the LLM can utilize its internally stored information more effectively, resulting in better performance on tasks requiring logical reasoning, multi-step computation, or decision-making.
@@ 3 Paradigms of Knowledge Injection
This section identifies four major paradigms of knowledge injection: Dynamic Knowledge Injection, Static Knowledge Embedding, Adapters, and Prompt Optimization. These paradigms illustrate various mechanisms by which external domain-specific knowledge can be incorporated into LLMs. We utilize unified notations, as described in Table 1, to systematically represent the processes. External knowledge $\mathcal{K}$ is integrated into LLMs either by modifying the original parameters $\theta$, introducing additional parameters $\phi$, or leveraging auxiliary mechanisms.
Figure 2: Four knowledge injection paradigms for LLMs. (a) Dynamic Knowledge Injection retrieves external knowledge during inference for enhanced reasoning. (b) Static Knowledge Injection embeds external knowledge into model parameters during fine-tuning. (c) Modular Knowledge Adapters use plug-and-play modules to dynamically adapt to tasks or updates. (d) Prompt Optimization utilizes precise prompts to guide the LLM without altering its parameters.
@@ 3.1 Dynamic Knowledge Injection
where $\mathbf{x}$ represents the original input, $\mathcal{R}$ denotes the retrieval function, $\mathcal{K}$ is the external knowledge base, and $\theta$ are the model parameters, which remain unchanged. This paradigm offers several advantages, including ease of updating (hence the term ”dynamic injection”) and the ability to incorporate new knowledge without retraining the model. However, it also presents challenges, such as dependency on the quality of the knowledge base $\mathcal{K}$, the retrieval function $\mathcal{R}$, and limitations imposed by the maximum input length of the LLM.
@@ 3.2 Static Knowledge Embedding
where $\mathcal{K}$ is the domain-specific knowledge base containing training samples $\mathbf{x_{s}}$ and $\mathbf{y_{s}}$, and $\mathcal{L}$ is a typical supervised training loss function. After optimization, the updated parameters $\Delta\theta$ are obtained.
At inference time, no further retrieval or external knowledge calls are required:
$\mathbf{y}=M\bigl{(}\mathbf{x};\Delta\theta\bigr{)}.$
This paradigm provides fast inference, as it eliminates additional retrieval steps and often results in stronger performance. However, it also comes with challenges, such as costly updates—requiring fine-tuning whenever domain knowledge changes—and scalability issues, as embedding large or frequently changing knowledge bases can be computationally expensive.
@@ 3.3 Modular Knowledge Adapters
where $M(\mathbf{x_{s}};\theta,\phi)$ represents the base model’s generation function enhanced with the new adapter parameters. At inference time, the enhanced model generates outputs as:
$\mathbf{y}=M\bigl{(}\mathbf{x};\theta,\phi\bigr{)}.$
This paradigm offers a parameter-efficient method to adapt LLMs to specific domains without modifying the original model weights. By freezing the base model’s parameters, the approach seeks to preserve previously acquired knowledge while enabling the seamless incorporation of new domain-specific information. However, this approach also introduces challenges, such as the need to design new architectural components and determine appropriate hyperparameters, including the size and number of adapters. These additional elements can increase the overall complexity of the model and its training process.
@@ 3.4 Prompt Optimization
Unlike previous approaches, prompt optimization does not retrieve knowledge from external sources. Instead, it focuses on fully leveraging or guiding the LLM to utilize its internal, pre-existing knowledge. The process can be formalized as:
$\mathbf{y}=M\bigl{(}[\mathbf{p},\mathbf{x}];\theta\bigr{)},$
where $\mathbf{p}$ represents a textual prompt containing implicit domain knowledge or specific instructions.
Prompt optimization offers significant advantages, including eliminating dependency on external domain knowledge bases and avoiding training. However, it also presents challenges, as designing effective prompts can be both complex and time-consuming. Additionally, long prompts may reduce the available context window, potentially affecting model efficiency and performance.
@@ 3.5 Comparison of the Four Paradigms
Table 2: Guidance on selecting knowledge injection paradigms based on training cost, inference speed, and limitations.
Dynamic knowledge injection integrates external knowledge at runtime, offering flexibility and adaptability to new information without increasing training cost. However, it requires an effective retrieval module, and the inference speed depends heavily on retrieval performance, which can slow down the overall process. Static knowledge embedding embeds domain expertise during pretraining or fine-tuning, requiring large-scale domain-specific data and significant training resources, including GPUs and time. While it incurs no additional inference cost, its limitations lie in the potential risks of catastrophic forgetting and its inability to adapt to evolving information. Modular adapters serve as a middle ground, allowing plug-and-play components to enhance domain-specific capabilities with minimal training data. Only a small subset of parameters needs to be trained, which reduces training costs, and inference speed is largely unaffected. However, the quality of training data significantly impacts the performance of this method. Prompt Optimization, on the other hand, avoids retraining entirely by leveraging carefully crafted inputs. It has no impact on inference speed but relies on extensive human effort to find optimal prompts. This method is limited in its ability to utilize new knowledge and primarily activates pre-existing knowledge.
We summarize these comparisons in Table 2 as a practical guide to help determine the most suitable method based on specific task requirements and scenarios.
@@ 4 Applications
4.1 Biomedicine
The biomedicine domain benefits from a wealth of specialized corpora, such as PubMed (Dernoncourt and Lee, 2017) and MedQA (Jin et al., 2021), enabling the development of LLMs specifically trained on biomedical texts. These models often follow the static knowledge embedding approach, leveraging the domain-specific richness of biomedical data. For instance, PMC-LLaMA (Wu et al., 2023) extends the LLaMA 7B model through further pretraining on 4.9 million PubMed Central articles curated from the S2ORC dataset (Lo et al., 2020), completing five epochs to embed biomedical knowledge effectively. Similarly, Med-PaLM 2 (Singhal et al., 2023b) builds on PaLM 2 via instruction fine-tuning. This fine-tuning incorporates a diverse mix of medical question-answering datasets, including MedQA, MedMCQA (Pal et al., 2022), and HealthSearchQA (Singhal et al., 2023a).
Beyond foundational models, integrating external tools and knowledge can further enhance performance. For example, GeneGPT (Jin et al., 2024) utilizes an LLM pretrained on code tasks to tackle GeneTuring tests by employing NCBI Web APIs. This approach combines in-context learning with an augmented decoding algorithm that identifies and executes API calls. Similarly, Med-PaLM (Singhal et al., 2023a) introduces vector prompts—representations that store and retrieve medical domain knowledge—to extend the capabilities of Flan-PaLM (Chung et al., 2024).
@@ 4.1 Biomedicine
Biomedicine PMC-LLaMA (Wu et al., 2023) Static Knowledge Embedding PMC-OA, MedC-I, PubMedQA, MedMCQA, USMLE Link
DALK  (Li and others, 2024)
Dynamic Knowledge Injection
Prompt Optimization
MedQA, MedMCQA, MMLU, QA4MRE Link
ChronicCareGPT (Liu et al., 2024b) Prompt Optimization eRisk Link
SA-MDKIF (Xu et al., 2024b) Modular Knowledge Adapters MedQuA,emrQA, PubMedQA, MedQA \
MaLP (Zhang et al., 2024b) Modular Knowledge Adapters HealthCareMagic-100k, iCliniq Link
BiomedRAG (Li et al., 2024) Dynamic Knowledge Injection CHEMPROT,DDI,ade-corpus-v2,MTsample,ADInt,UMLS Link
MedINST (Han et al., 2024) Static Knowledge Embedding MedINST Link
Finance FLANG  (Shah and others, 2022) Static Knowledge Embedding
@@
Fin-LLaMA  (Konstantinidis and others, 2024) Static Knowledge Embedding fin-llama-dataset Link
@@
Google API, Docker container, Internet,
Hardware API document, Physical world hardware
@@
FUSE (Liu et al., 2024f) Prompt Optimization True News Dataset, Big Five Personality Traits \
@@
4.2 Finance
Fine-tuned financial LLMs have demonstrated significant advancements in adapting general-purpose models to domain-specific tasks through task-specific training. PIXIU (Xie et al., 2023) fine-tunes LLaMA on 136K instruction samples tailored to financial tasks, enabling the model to handle a wide range of domain-relevant scenarios. Instruct-FinGPT (Zhang et al., 2023) fine-tunes LLaMA on 10K instruction samples derived from two financial sentiment analysis datasets, focusing primarily on finance classification tasks. FinGPT (Yang et al., 2023) introduces a comprehensive end-to-end framework for training and deploying FinLLMs in the financial industry. Utilizing the LoRA technique, FinGPT fine-tunes open-source LLMs like LLaMA and ChatGLM with approximately 50K task-specific samples, achieving efficient fine-tuning without full model retraining.
In contrast, scratch-trained financial LLMs aim to create models specifically designed for financial tasks from the ground up. BloombergGPT (Wu and others, 2023) leverages a subset of 5 billion tokens from Bloomberg-specific data, representing only 0.7% of its total training corpus, to tailor its model to financial applications. XuanYuan 2.0 (Zhang and Yang, 2023) combines 366 billion tokens for pre-training with an additional 13 billion tokens for fine-tuning, creating the largest Chinese financial chat model. Similarly, Fin-T5 (Lu et al., 2023) introduces a Chinese financial pre-training language model built on the T5 architecture, utilizing a 300GB financial corpus. Furthermore, SNFinLLM (Zhao et al., 2024a) dynamically incorporates real-time financial data during inference to enhance decision-making capabilities, demonstrating the value of domain-specific pretraining and adaptability in financial LLMs.

### v2

@@ Abstract
Injecting Domain-Specific Knowledge into Large Language Models:
A Comprehensive Survey
Zirui Song1∗   Bin Yan2thanks: Equal contributions.   Yuhan Liu3   Miao Fang2   Mingzhe Li4   Rui Yan5   Xiuying Chen1thanks: Corresponding author. 1Mohamed bin Zayed University of Artificial Intelligence       2Northeastern University       3Gaoling School of Artificial Intelligence, Renmin University of China       4ByteDance       5Wuhan University E-mail: zirui.song@mbzuai.ac.ae, xiuying.chen@mbzuai.ac.ae
@@ 1 Introduction
LLMs have achieved extraordinary success across various tasks, showcasing remarkable capabilities in reasoning, knowledge representation, and decision-making Song et al. (2025); Xu et al. (2025). However, despite their impressive performance in general-purpose applications, many specialized domains, such as healthcare, chemistry, and legal analysis, demand the integration of domain-specific knowledge to achieve high accuracy and reliability. To address this challenge, researchers have explored methods to enhance LLMs through external or embedded domain expertise, a process often referred to as knowledge injection, as shown in Figure 1. This approach aims to bridge the gap between general-purpose language understanding and the stringent requirements of domain-specific tasks, enabling LLMs to perform effectively in highly specialized contexts.
Figure 1: Illustration of Growth Trends in Domain-Specific Knowledge Injection into LLMs. The chart displays the cumulative number of papers published between October 2022 and December 2024. Different colors and border styles represent various injection methods and domains.
Building on the foundational capabilities of general-purpose LLMs, knowledge injection techniques provide an effective means to address their limitations in handling specialized applications. Compared to the generalized approach of standard LLMs, knowledge injection offers two key advantages: 1) incorporating precise, domain-specific knowledge to improve accuracy and reliability in specialized tasks, and 2) allowing LLMs to dynamically adapt to new information or evolving knowledge bases, ensuring up-to-date expertise. These techniques bridge the gap between general-purpose understanding and domain-specific demands by leveraging both structured and unstructured knowledge sources. As a result, knowledge injection methods have been successfully applied in fields such as healthcare, chemistry, and legal analysis, significantly enhancing LLM performance. For example, biomedical LLMs  (Cho and Lee, 2025; Bolton et al., 2024; Yan et al., 2023) have demonstrated superior accuracy in tasks like medical diagnostics and regulatory compliance, while domain-specific models for material science (Tang et al., 2025a; Xie et al., 2024; Antunes et al., 2024; Zhang et al., 2024b) have achieved advances in material property prediction and discovery. These dedicated models underscore the transformative potential of integrating domain knowledge into LLMs.
@@
We begin in Section 2 with background on domain-specific knowledge and its role in LLMs. Section 3 presents a unified framework of four knowledge injection paradigms: (1) Dynamic Knowledge Injection at inference time; (2) Static Knowledge Embedding during training or fine-tuning; (3) Modular Adapters for parameter-efficient integration; and (4) Prompt Optimization via carefully designed inputs. Section 4 examines these methods across domains such as materials science, chemistry, biology, and law. Section 5 summarizes key datasets, tools, and comparative results. Section B outlines open challenges, including scalability, robustness, and domain transfer. Finally, Section 7 concludes the paper and reflects on future directions.
@@ 2.1 Domain-Specific Knowledge
Domain-specific knowledge refers to specialized information or expertise pertinent to a specific field or application, distinguishing it from general knowledge that spans across multiple domains. While general knowledge enables models to understand broad contexts, domain-specific knowledge is essential for addressing specialized tasks where precise, field-specific understanding is required Wang et al. (2025); Li et al. (2025a); Zhang et al. (2025). For instance, in scientific text processing (Bran et al., 2023), models must comprehend complex scientific terminologies, concepts, and methodologies to provide accurate and relevant answers. In e-commerce search (Zhao et al., 2024a), understanding domain-specific terms such as product categories, technical specifications, or colloquial shopping language is crucial for delivering relevant search results and recommendations. In healthcare applications, LLMs must understand medical terminologies, diagnoses, treatment plans, and drug interactions. For example, biomedical question answering (Singhal et al., 2025; Pei et al., 2024) and medical report summarization rely on integrating knowledge from medical literature like PubMed (Dernoncourt and Lee, 2017). To address these needs, researchers have explored various methods for incorporating domain-specific knowledge into LLMs. In this paper, we aim to provide a survey of these various injection methods.
@@ 2.2 Knowledge Representation and Encoding
Knowledge can take different forms depending on structure and application needs. Knowledge graphs (Liao et al., 2025; Zhang et al., 2024d) encode entities and their relationships in a structured graph, supporting reasoning and inference in tasks like question answering. In contrast, text-based sources like Wikipedia (Jeong et al., 2024) provide rich but unstructured information, useful for tasks requiring broad contextual understanding. Knowledge can also be stored in vector space rather than in text or graph form. For example, soft prompt tuning (Peng et al., 2025; Singhal et al., 2023a) embeds useful information as vectors, which are appended to inputs to guide LLMs on specific tasks. Beyond external forms, knowledge may also emerge internally: chain-of-thought prompting (Sanwal, 2025; Yao et al., 2024) introduces intermediate reasoning steps that help LLMs decompose complex problems and access internal knowledge more effectively—improving performance in tasks involving reasoning, multi-step computation, or decision-making.
2.3 Knowledge Injection Survey
Prior surveys on knowledge-enhanced language models vary in focus and scope. The most relevant works include the following: Cadeddu et al. (2024), who focus on scientific article classification and offer practical insights but lack broader methodological generalization; Wang et al. (2024b), who focus on knowledge editing and aim to update internal model knowledge with minimal side effects; and Hu et al. (2023), who adopt a model-centric perspective by classifying knowledge-enhanced models based on task type and knowledge source, though they primarily cover pre-LLM architectures such as BERT and ERNIE. In contrast, our work presents a unified view of knowledge injection in LLMs, emphasizing capability enhancement through external knowledge integration across diverse tasks.
@@ 3 Paradigms of Knowledge Injection
To systematically understand how domain knowledge is integrated into LLMs, we categorize existing approaches into four paradigms based on when the knowledge is incorporated and how it interacts with the model, as shown in Figure 2. Specifically, Static Knowledge Injection and Modular Knowledge Adapters integrate knowledge prior to inference and involve parameter updates—through either full fine-tuning or adapter-based tuning. In contrast, Dynamic Knowledge Injection and Prompt Optimization inject knowledge at inference time without altering model parameters: the former retrieves external information, while the latter leverages internal knowledge through designed prompts.
We utilize unified notations, as described in Table 1, to systematically represent the processes.
Figure 2: Four knowledge injection paradigms for LLMs. (a) Dynamic Knowledge Injection retrieves external knowledge during inference. (b) Static Knowledge Injection embeds external knowledge into model parameters during fine-tuning. (c) Modular Knowledge Adapters use plug-and-play modules to dynamically adapt to tasks. (d) Prompt Optimization utilizes precise prompts to guide the LLM without altering its parameters.
@@ 3.1 Dynamic Knowledge Injection
where $\mathbf{x}$ represents the original input, $\mathcal{R}$ denotes the retrieval function, $\mathcal{K}$ is the external knowledge base, and $\theta$ are the model parameters, which remain unchanged. This paradigm offers several advantages, including ease of updating (hence the term "dynamic injection") and the ability to incorporate new knowledge without retraining the model. However, it also presents challenges, such as dependency on the quality of the knowledge base $\mathcal{K}$, the retrieval function $\mathcal{R}$, and limitations imposed by the maximum input length of the LLM. To improve retrieval quality, commonly used techniques include semantic matching based on sentence embeddings and efficient vector databases for fast similarity search.
@@ 3.2 Static Knowledge Embedding
where $\mathcal{K}$ is the domain-specific knowledge base containing training samples $\mathbf{x_{s}}$ and $\mathbf{y_{s}}$, and $\mathcal{L}$ is a training loss function. After optimization, the updated parameters $\Delta\theta$ are obtained.
At inference time, no further retrieval or external knowledge calls are required: $\mathbf{y}=M\bigl{(}\mathbf{x};\Delta\theta\bigr{)}$. This paradigm enables fast inference by removing the need for additional retrieval steps and often delivers stronger performance. However, it also presents challenges, such as high update costs since fine-tuning is required when domain knowledge changes, and scalability concerns because embedding large or frequently changing knowledge bases demands significant computational resources.
@@ 3.3 Modular Knowledge Adapters
where $M(\mathbf{x_{s}};\theta,\phi)$ represents the base model’s generation function enhanced with the new adapter parameters. At inference time, the enhanced model generates outputs as: $\mathbf{y}=M\bigl{(}\mathbf{x};\theta,\phi\bigr{)}$. This paradigm offers a parameter-efficient method to adapt LLMs to specific domains without modifying the original model weights. By freezing the base model’s parameters, the approach seeks to preserve previously acquired knowledge while enabling the seamless incorporation of new domain-specific information. However, this approach also introduces challenges, such as the need to design new architectural components and determine appropriate hyperparameters, including the size and number of adapters. These additional elements can increase the overall complexity of the model and its training process.
@@ 3.4 Prompt Optimization
Prompt optimization refers to the practice of guiding LLMs to perform domain-specific tasks by crafting effective textual prompts. Unlike retrieval-based methods, it relies entirely on the model’s internal knowledge and does not require access to external knowledge bases or fine-tuning. The process can be formalized as:
$\mathbf{p}^{*}=\textstyle\arg\min_{\mathbf{p}}\mathcal{L}\big{(}M([\mathbf{p},\mathbf{x}];\theta),\mathbf{y}^{*}\big{)},$
where $\mathbf{p}$ is a prompt containing domain-relevant cues, $\mathbf{x}$ is the task input, and $\theta$ are the fixed parameters of the LLM.
This paradigm offers practical advantages such as lightweight deployment, no training overhead, and adaptability across domains. However, it also faces key challenges: designing prompts that elicit accurate responses can be non-trivial, and long prompts may reduce the available input space due to context length limitations. Prompt-based approaches can be broadly categorized into manual prompting, prompt tuning, and prefix tuning. These differ in how prompts are constructed or optimized—ranging from discrete, static prompts to learnable embeddings—and have been widely adopted for low-resource domain adaptation.
@@ 3.5 Comparison of the Four Paradigms
Table 2: Guidance on choosing injection methods by training cost, inference speed, and constraints.
Dynamic knowledge injection introduces external knowledge at runtime, offering flexibility and adaptability without added training cost. However, it depends on efficient retrieval, and inference speed can suffer if retrieval performance is poor. Static knowledge embedding integrates domain expertise during pretraining or fine-tuning, requiring large-scale data and significant computational resources. It adds no inference cost but struggles to adapt to new information and is prone to catastrophic forgetting. Modular adapters offer a middle ground by enhancing domain capabilities through plug-and-play modules that require minimal training data. Only a small number of parameters are trained, reducing cost and preserving inference speed, though performance heavily depends on data quality. Prompt optimization avoids retraining by using well-crafted inputs. It maintains fast inference but relies on significant manual effort and is limited to activating existing knowledge rather than incorporating new information. We summarize these comparisons in Table 2 as a practical guide to help determine the suitable method based on specific requirements and scenarios.
@@ 4 Applications
4.1 Finance
In the financial domain, LLM generally follow two main development paths: fine-tuning general-purpose models on financial tasks or training models from scratch using domain-specific corpora.
For fine-tuning, PIXIU(Xie et al., 2023) adapts LLaMA using 136K financial instruction samples, equipping the model to handle diverse finance-related scenarios. Instruct-FinGPT(Zhang et al., 2023) focuses on sentiment classification by fine-tuning on 10K samples from two financial sentiment datasets. FinGPT (Yang et al., 2023) proposes an end-to-end framework for developing FinLLMs, efficiently fine-tuning LLaMA and ChatGLM with 50K samples via LoRA, significantly reducing computational costs. In contrast, scratch-trained FinLLMs aim for deep domain alignment from the ground up. BloombergGPT (Wu et al., 2023b) uses 5B Bloomberg-specific tokens (0.7% of its total corpus) to specialize in financial applications. XuanYuan 2.0 (Zhang and Yang, 2023) is the largest Chinese financial chatbot, trained on 366B tokens and fine-tuned on 13B. Fin-T5 (Lu et al., 2023) leverages a 300GB Chinese financial corpus using the T5 architecture, while SNFinLLM (Zhao et al., 2024a) enhances inference through real-time financial data injection.
In summary, this field showcases a rich diversity in training strategies, from lightweight tuning to comprehensive, end-to-end development.
@@ 4.1 Finance
Finance FLANG  (Shah et al., 2022) Static Knowledge Embedding
@@
Fin-LLaMA  (Konstantinidis et al., 2024) Static Knowledge Embedding fin-llama-dataset Link
@@
Fino1  (Qian et al., 2025) Static Knowledge Embedding
FinQA, TATQA , DocMath (Simpshort and Compshort),
DocFinQA , and BizBench-QA
Link
Biomedicine PMC-LLaMA (Wu et al., 2023a) Static Knowledge Embedding PMC-OA, MedC-I, PubMedQA, MedMCQA, USMLE Link
DALK  (Li et al., 2024a)
Dynamic Knowledge Injection
Prompt Optimization
MedQA, MedMCQA, MMLU, QA4MRE Link
ChronicCareGPT (Liu et al., 2024b) Prompt Optimization eRisk Link
SA-MDKIF (Xu et al., 2024c) Modular Knowledge Adapters MedQuA,emrQA, PubMedQA, MedQA \
MaLP (Zhang et al., 2024c) Modular Knowledge Adapters HealthCareMagic-100k, iCliniq Link
BiomedRAG (Li et al., 2024b) Dynamic Knowledge Injection CHEMPROT,DDI,ade-corpus-v2,MTsample,ADInt,UMLS Link
MedINST (Han et al., 2024) Static Knowledge Embedding MedINST Link
K-COMP (Cho and Lee, 2025) Dynamic Knowledge Injection MedCorp corpus \
OntoTune (Liu et al., 2025c) Static Knowledge Embedding SemEval2018 Task 9 dataset Link
@@
college chemistry textbooks:Quantum chemistry (quan), hemistry kinetics (matter)
Quantum mechanics (chemmc), Physical chemistry (atkins)
@@
LLaMat  (Mishra et al., 2025) Static Knowledge Embedding MatBookQA,MaScQA,MatSciInstruct \
OmniScience (Prabhakar et al., 2025) Static Knowledge Embedding daring-anteater dataset, s1K dataset \
@@
FUSE (Liu et al., 2025b) Prompt Optimization True News Dataset, Big Five Personality Traits \
@@
4.2 Biomedicine
The biomedicine domain benefits from a wealth of specialized corpora, such as PubMed (Dernoncourt and Lee, 2017) and MedQA (Jin et al., 2021), enabling the development of LLMs specifically trained on biomedical texts. These models often follow the static knowledge embedding approach, leveraging the domain-specific richness of biomedical data. For instance, PMC-LLaMA (Wu et al., 2023a) extends the LLaMA 7B model through further pretraining on 4.9 million PubMed Central articles curated from the S2ORC dataset (Lo et al., 2020), completing five epochs to embed biomedical knowledge effectively. Similarly, Med-PaLM 2 (Singhal et al., 2023b) builds on PaLM 2 via instruction fine-tuning. This fine-tuning incorporates a diverse mix of medical question-answering datasets, including MedQA, MedMCQA (Pal et al., 2022), and HealthSearchQA (Singhal et al., 2023a).
Beyond foundational models, integrating external tools and knowledge can further enhance performance. GeneGPT(Jin et al., 2024) leverages a code-pretrained LLM to address GeneTuring tests by calling NCBI Web APIs, combining in-context learning with an augmented decoding algorithm capable of identifying and executing API requests. Med-PaLM(Singhal et al., 2023a) extends the capabilities of Flan-PaLM (Chung et al., 2024) through the use of vector prompts—dense representations designed to store and retrieve medical domain knowledge during inference.
Overall, biomedical LLMs lead in combining static pretraining, instruction tuning, and tool integration, reflecting a shift toward hybrid reasoning in specialized AI.